# IceCube Collaboration

### v1

@@ Abstract
We report a measurement of the inelasticity distribution in the scattering of neutrinos of energy $80-560~{}\mathrm{GeV}$ off nucleons, which is sensitive to the inclusive differential cross section. This analysis is based on a sample of atmospheric muon neutrinos detected in the IceCube sub-array DeepCore during 2012-2021, and is the first such measurement in this energy range. Our measurement extends to energies where accelerator data is not available, hence we compare our results to predictions from perturbative QCD calculations, finding good agreement.
footnotetext: email: analysis@icecube.wisc.edu (corresponding author)
@@ I Introduction
Another source of information on neutrino interactions are neutrino telescopes which detect naturally created neutrinos. These experiments access much higher energies, but are limited by the uncertainties on the incoming flux and the accuracy with which an interaction can be reconstructed. Despite these limitations, these experiments can contribute in a complimentary way to the understanding of neutrino cross sections. IceCube, which detects both atmospheric and astrophysical neutrinos, has previously reported measurements of the total neutrino cross section up to a PeV [5, 6], as well as a measurement of the inelasticity distribution at neutrino energies of $1-20$ TeV [7]. Such energies are well above these accessible at terrestrial accelerators.
Here, we present the results from a measurement of the flux-averaged inelasticity distribution in neutrino-nucleon interactions from 80 GeV to 560 GeV using data from the IceCube DeepCore array. At these energies neutrinos interact mainly through deep-inelastic scattering (DIS), which involves scattering off a quark or gluon in the nucleon, resulting in a lepton and a hadronic shower in the final state. The inelasticity $y$ is a measure of the energy distribution of the products of the interaction, defined as [8]
@@
A measurement of the inelasticity of neutrino-nucleon interactions simultaneously probes the fundamental properties of the weak force and our understanding of the structure of the nucleon in QCD. At energies below 1 PeV neutrinos and antineutrinos have different inelasticity distributions as the bulk of their interactions happen on valence quarks. The difference disappears at high energies as neutrinos and antineutrinos are able to interact with equal numbers of sea quarks and anti-quarks and with gluons [8]. The measurement presented here uses the atmospheric neutrino flux, which contains both neutrinos and antineutrinos. We measure a convolution of the neutrino and antineutrino inelasticity distributions with the neutrino to antineutrino ratio. Such a measurement can validate current cross section and flux models, while any deviation in the measured inelasticity distribution could indicate a mismodeling either in cross section or in atmospheric flux predictions. We also invert our results to measure the neutrino fraction in our sample, assuming a specific cross section calculation.
@@
At $E_{\nu}=$ 80-560 GeV, neutrino-nucleon interactions are expected to be dominated by DIS, however, resonance (RES) and quasi-elastic (QE) scattering events provide a non-negligible contribution at low inelasticity values. In this study we measure the inclusive inelasticity distribution. This is because in DeepCore we cannot reliably differentiate between QE, RES and DIS events due to the large inter-module spacing compared to the characteristic size of the hadronic cascades.
@@
We begin by introducing the IceCube Neutrino Observatory (Sec. II), followed by a description of the data sample used (Sec. III). The analysis is presented in Sec. IV, and results in Sec. V. In Sec. VI we discuss the implications of our results. The limitations of the parametrization used in our analysis are described in detail in Appendix A. In Appendix B we show a detailed comparison of the impact of measured and nuisance parameters on our data sample.
II The IceCube neutrino detector
@@
The IceCube Neutrino Observatory is an ice-Cherenkov detector consisting of 5,160 Digital Optical Modules (DOMs) deployed deep under the surface at the geographic South Pole. The DOMs instrument a volume of approximately 1 km3, starting at a depth of 1,450 m below the surface and ending at 2,450 m. The bulk of the array is optimized for the observation of neutrinos with energies of 1 TeV and above, and has DOMs spaced by 17 m in the vertical direction, and 125 m in the horizontal direction [9]. A low-energy extension, known as DeepCore, is located at the bottom, central region of the detector, where higher quantum efficiency PMTs are deployed closer together, with a vertical separation of 7 m and a horizontal separation between 40 m and 70 m [10]. DeepCore can observe neutrino events down to $\sim$10 GeV [11].
@@
The sensitive element in the DOMs is a 10-inch photomultiplier tube (PMT) facing downward away from the ice surface [12]. The PMT is enclosed in a glass sphere together with the electronics for power and signal digitization [9]. The waveforms produced by the PMTs in each DOM are recorded by analog-digital converters and the digitized data is passed to a pulse-finding algorithm, that deconvolves it into pulses using the typical response of the PMT to single photons [13, 14]. The interpretation of the properties of the events, such as the event selection and reconstruction described below, is carried out based on the information contained in these pulses.
@@
In this analysis we use DeepCore to study the inelasticity of events with primary neutrino energies in the range from $80$ to $560$ GeV. The hadronic cascade in such events are still mostly contained inside the more densely instrumented DeepCore volume, which gives us a more precise reconstruction compared to the main IceCube array. The energy of the muons exiting the DeepCore, but still contained inside the IceCube, can be estimated from their track length.
@@ III Data selection
The data sample used for this analysis comes from the common DeepCore event selection described in Ref. [11], specifically at Level 5. The common selection focuses on rejecting events produced by atmospheric muons as well as detector noise, while keeping events where the neutrino interacted inside the DeepCore region. The sample after the common selection has approximately a $2:1$ ratio of neutrinos to non-neutrino backgrounds, so further levels of selection described in this section, are necessary for refining a high purity $\nu_{\mu}$ CC-only sample for this study.
@@ III.1 Event reconstruction and classification
Table 1: Sample composition at the final level of selection, as described in Sec. III, with projected event counts for 9.28 years of data, calculated using the baseline Monte-Carlo set.
@@
The second BDT classifier determines the Particle Identification Score (PID), which divides events into track-like and cascade-like. The classifier is trained to identify $\nu_{\mu}$ CC interactions as track-like events, while $\nu_{e}$ CC interactions represent cascade-like events. The BDT uses input of five variables coming from the reconstruction algorithm: track length of the outgoing muon, cascade energy, zenith angle, zenith angle uncertainty, and the likelihood difference between the best fit hypothesis with arbitrarily long track and with track length fixed at zero. A comparison of PID Classifier score for data and simulation is shown on Fig. 2, which also indicates the cut on PID score $>$ 0.7 described in Sec. III.3.
Figure 1: Muon Classifier score comparison for data and for predicted sample composition at the reconstruction level. The Monte-Carlo normalization is scaled to match total event counts to data. The grey dashed line indicates the position of the cut on Muon Classifier score, with values to the left of the line removed from the final sample.
Figure 2: PID score comparison for data and for predicted sample composition at the reconstruction level. The Monte-Carlo normalization is scaled to match total event counts to data. The grey dashed line indicates the position of the cut on PID score, with values to the left of the line removed from the final sample.
Figure 3: Level 4 Noise Classifier score comparison for data and for predicted sample composition at the reconstruction level. The Monte-Carlo normalization is scaled to match total event counts to data. The grey dashed line indicates the position of the additional cut on the classifier score, with values to the left of the line removed from the final sample.
@@
In order to further reduce non-neutrino background, the Muon Classifier score is required to be $>0.4$ for events to pass to the next step of selection. An additional cut of $>0.95$ is applied on the Level 4 Noise Classifier [11] to reduce remaining background. The performance of the Level 4 Noise Classifier score for our sample at the reconstruction level is shown in Figure 3.
@@ III.2 Rejection of non-neutrino backgrounds
Another class of variables used to remove noise and low quality events are calculated based on estimated photon speeds, under the assumption that each detected photon has started at the event interaction vertex and and then travelled without scattering to the corresponding DOM. Photon speeds estimated in this way will not necessarily accurately represent the real event, however the variables calculated using these are still useful, as they provide information on the event quality. In our event selection, we require the median estimated speed of all detected photons in an event to be $>0.4c$, where $c$ is the speed of light in vacuum. We also require that in each event the fraction of detected photons with estimated speeds below zero or above $c$ is $<0.4$, because such pulses typically result from the PMT noise. A photon speed below zero indicates that the pulse happened before the reconstructed interaction time, while photon speed above $c$ means that the pulse happened either too far from the reconstructed vertex or too early to be explained by a photon travelling from the vertex at the speed $c$. While a photon travelling in ice will have a speed of $c/n$, where $n$ is the refractive index of ice, the choice was made nevertheless to make a cut at the speed of $c$, to avoid penalizing events where a muon travels at speed higher then $c/n$ and than produces a Cherenkov photon near the DOM.
Restrictions are also applied to reject neutrino events with likely coincident atmospheric muons entering the detector. This is achieved by introducing cuts on events with pulses in the veto region. This region includes the top and the outermost optical modules of IceCube array, as these are more likely to detect light from atmospheric muons entering the detector. We limit the allowed number of pulses in the outermost layer of IceCube to be below 8. At the same time we also require that there is no downward travelling trend in the pulse distribution in the top 15 layers of DOMs in IceCube, which corresponds to a vertical range of roughly 255 m.
We introduce vertex containment cuts on reconstructed vertex position to remove events with vertex located outside the DeepCore. We require that the vertical position of the vertex be in the range -500 m $<z_{\mathrm{reco}}<$200 m from the vertical center of IceCube and that horizontal radial distance from the central string be $\rho_{\mathrm{reco}}<$ 300 m.
@@ III.3 Rejection of non-$\nu_{\mu}$ CC backgrounds
The remaining steps focus on selecting $\nu_{\mu}$ CC events with high quality reconstruction. The primary event types targeted for the removal are very dim atmospheric muons, which leave few or no pulses on the veto region, and were therefore missed by previous selection steps, and non-$\nu_{\mu}$ CC neutrino interactions.
@@ III.4 Final level sample
The composition of the final sample is given in Table 1. Our selection results in a signal purity of 99%. The average ratio of neutrinos to antineutrinos is 2.4, which reflects the properties of the atmospheric neutrino flux, the neutrino cross section, as well as event selection efficiency dependence on true inelasticity. The data does not provide sufficient information to distinguish between neutrinos and antineutrinos, so the analysis is done for the sum of both.
The reconstruction resolutions for neutrino energy and inelasticity in the final simulated sample are presented in Fig. 4. The energy estimator shows a clear correlation with neutrino energy up to 500 GeV. Above that, the resolution degrades, because the typical size of the hadronic cascade approaches the size of DeepCore and a larger portion of the muon track could be located outside DeepCore. The inelasticity estimator also displays a correlation across all values, with deviations observed towards the edges, as is expected from one-sided errors in those regions. These two reconstructed quantities are used to extract the inelasticity measurement, as described in the next section.
@@ IV.1 General strategy
Figure 7: Fit of the parametrization (3) to inclusive inelasticity distributions for generator level GENIE Monte-Carlo. Fits were done separately for neutrino and antineutrino inclusive inelasticity distributions in three energy bins, which were weighted to reflect energy distributions in energy bins at analysis level. Fit results for parameters $\langle y\rangle$ and $\log_{10}\lambda$ are shown in the bottom left of each plot.
Figure 8: Data/Monte-Carlo agreement post-fit for the reconstructed inelasticity distribution in each of three $E_{\mathrm{reco}}$ bins.
This analysis is performed by passing data and simulation through the same selection and reconstruction steps, and then comparing their distributions in the space of observables chosen. The shape of the true inelasticity distribution of the simulation is parametrized, so it can be modified, along with nuisance parameters that describe the impact of different sources of uncertainties, in order to find the values that result in the best fit to the data. In this study, we compare a two-dimensional histogram in $E_{\mathrm{reco}}$ and $y_{\mathrm{reco}}$ with 3 bins in reconstructed energy between 100 GeV and 1 TeV and 11 equally sized bins in reconstructed inelasticity, between zero and one.
@@ IV.2 Parametrization of inelasticity
To parameterize the shape of the inelasticity distribution, our study takes advantage of the expected dependence of the neutrino-nucleon interaction cross section on inelasticity, as described in Sec. I. The functional form, used in previous IceCube studies [7, 19, 20], is the probability distribution of inelasticity and depends on two parameters $\epsilon$ and $\lambda$,
@@
where $\sigma$ is total cross section for a given energy, $\frac{d\sigma}{dy}$ is the differential cross section as function of inelasticity for a given energy and $N$ is a normalization factor. In order to treat the distribution as a probability density function it is normalized to one, so
@@
It is important to emphasize that this parametrization is an effective description of the range of “reasonable” inelasticity distribution shapes, however it has a number of limitations described in Appendix A.
The parametrization can be applied to both neutrino and antineutrino inelasticity distributions separately. Figure 7 shows Eq. 3 fitted to muon neutrino and antineutrino expectation from GENIE for three energy bins. However, due to our inability to differentiate between neutrino and antineutrino events, we use the following parametrization to describe the flux-averaged inelasticity distribution at energy $E_{\nu}$,
@@
where $\widetilde{\Phi}_{\nu(\bar{\nu})}$ is an energy-dependent fraction of neutrinos (antineutrinos) in our sample, which depends on atmospheric neutrino (antineutrino) flux $\Phi_{\nu(\bar{\nu})}$, the total neutrino(antineutrino)-nucleon cross section $\sigma_{\nu(\bar{\nu})}$, as well as event selection, which is indicated as $\mathrm{sel.}$ in the equation. $\frac{dp}{dy}_{\nu(\bar{\nu})}$ denotes individual neutrino (antineutrino) inelasticity distributions. It is important to highlight that in Eq. 6 $E_{\nu}$ refers to both neutrino and antineutrino energy.

### v2

# Measurement of the inelasticity distribution of neutrino-nucleon interactions for $\mathbf{80~{}GeV<E_{\nu}<560~{}GeV}$ with IceCube DeepCore

@@ Abstract
We report a study of the inelasticity distribution in the scattering of neutrinos of energy $80-560~{}\mathrm{GeV}$ off nucleons. Using atmospheric muon neutrinos detected in IceCube’s sub-array DeepCore during 2012–2021, we fit the observed inelasticity in the data to a parameterized expectation and extract the values that describe it best. Finally, we compare the results to predictions from various combinations of perturbative QCD calculations and atmospheric neutrino flux models.
DOI: 10.1103/PhysRevD.111.112001
footnotetext:
Published by the American Physical Society under the terms of the Creative Commons Attribution 4.0 International license. Further distribution of this work must maintain attribution to the author(s) and the published article’s title, journal citation, and DOI. Funded by SCOAP3.
@@ I Introduction
Another source of information on neutrino interactions are neutrino telescopes which detect naturally created neutrinos. These experiments access much higher energies, but are limited by the uncertainties on the incoming flux and the accuracy with which an interaction can be reconstructed. Despite these limitations, these experiments can contribute in a complimentary way to the understanding of neutrino cross sections. IceCube, which detects both atmospheric and astrophysical neutrinos, has previously reported measurements of the total neutrino cross section up to a PeV [5, 6] as well as a measurement of the inelasticity distribution at neutrino energies of $1-20$ TeV [7]. Such energies are well above those accessible at terrestrial accelerators.
Here, we present the results from a measurement of the flux-averaged inelasticity distribution in neutrino-nucleon interactions from 80 to 560 GeV using data from the IceCube DeepCore array. At these energies neutrinos interact mainly through deep-inelastic scattering (DIS), which involves scattering off a quark or gluon in the nucleon, resulting in a lepton and a hadronic shower in the final state. The inelasticity $y$ is a measure of the energy distribution of the products of the interaction, defined as [8]
@@
The inelasticity distribution of neutrino-nucleon interactions simultaneously probes the fundamental properties of the weak force and our understanding of the structure of the nucleon in QCD. At energies below 1 PeV neutrinos and antineutrinos have different inelasticity distributions as the bulk of their interactions happen on valence quarks. The difference disappears at high energies as neutrinos and antineutrinos are able to interact with equal numbers of sea quarks and antiquarks and with gluons [8]. The result presented here uses the atmospheric neutrino flux, which contains both neutrinos and antineutrinos. We measure a convolution of the neutrino and antineutrino inelasticity distributions with the neutrino to antineutrino ratio. Such a study can validate current cross section and flux models, while any deviation in the extracted inelasticity distribution could indicate a mismodeling either in cross section or in atmospheric flux predictions. We also invert our results to extract the neutrino fraction in our sample, assuming a specific cross section calculation.
@@
At $E_{\nu}=$ 80–560 GeV, neutrino-nucleon interactions are expected to be dominated by DIS; however, resonance (RES) and quasielastic (QE) scattering events provide a non-negligible contribution at low inelasticity values. In this study, we report the inclusive inelasticity distribution. This is because in DeepCore we cannot reliably differentiate between QE, RES, and DIS events due to the large intermodule spacing compared to the characteristic size of the hadronic cascades.
@@
We begin by introducing the IceCube Neutrino Observatory (Sec. II), followed by a description of the data sample used (Sec. III). The analysis is presented in Sec. IV, and results are presented in Sec. V. In Sec. VI, we discuss the implications of our results. The limitations of the parametrization used in our analysis are described in detail in Appendix A. In Appendix B, we show a detailed comparison of the impact of measured and nuisance parameters on our data sample.
II IceCube neutrino detector
@@
The IceCube Neutrino Observatory is an ice-Cherenkov detector consisting of 5160 digital optical modules (DOMs) deployed deep under the surface at the geographic South Pole. The DOMs instrument a volume of approximately 1 km3, starting at a depth of 1450 m below the surface and ending at 2450 m. The bulk of the array is optimized for the observation of neutrinos with energies of 1 TeV and above and has DOMs spaced by 17 m in the vertical direction and 125 m in the horizontal direction [9]. A low-energy extension, known as DeepCore, is located at the bottom, central region of the detector, where higher quantum efficiency photomultiplier tubes (PMTs) are deployed closer together, with a vertical separation of 7 m and a horizontal separation between 40 and 70 m [10]. DeepCore can observe neutrino events down to $\sim$10 GeV [11].
@@
The sensitive element in the DOMs is a 10 in. PMT facing downward away from the ice surface [12]. The PMT is enclosed in a glass sphere together with the electronics for power and signal digitization [9]. The waveforms produced by the PMTs in each DOM are recorded by analog-digital converters, and the digitized data are passed to a pulse-finding algorithm, which deconvolves it into pulses using the typical response of the PMT to single photons [13, 14]. The interpretation of the properties of the events, such as the event selection and reconstruction described below, is carried out based on the information contained in these pulses.
@@
In this analysis, we use DeepCore to study the inelasticity of events with primary neutrino energies in the range from $80$ to $560$ GeV. The hadronic cascade in such events is still mostly contained inside the more densely instrumented DeepCore volume, which gives us a more precise reconstruction compared to the main IceCube array. The energy of the muons exiting the DeepCore, but still contained inside the IceCube, can be estimated from their track length.
@@ III Data selection
The data sample used for this analysis comes from the common DeepCore event selection described in Ref. [11], specifically at Level 5. The common selection focuses on rejecting events produced by atmospheric muons as well as detector noise, while keeping events where the neutrino interacted inside the DeepCore region. Based on Monte Carlo modeling, the sample has approximately a $2:1$ ratio of neutrinos to non-neutrino backgrounds after the common selection, so further levels of selection described in this section are necessary for refining a high purity $\nu_{\mu}$ CC-only sample for this study.
@@ III.1 Event reconstruction and classification
Table 1: Sample composition at the final level of selection, as described in Sec. III, with projected event counts for 9.28 years of data, calculated using the baseline Monte Carlo set.
@@
The second BDT classifier determines the Particle Identification Score (PID), which divides events into tracklike and cascadelike. The classifier is trained to identify $\nu_{\mu}$ CC interactions as tracklike events, while $\nu_{e}$ CC interactions represent cascadelike events. The BDT uses input of five variables coming from the reconstruction algorithm: track length of the outgoing muon, cascade energy, zenith angle, zenith angle uncertainty, and the likelihood difference between the best-fit hypothesis with arbitrarily long track and with track length fixed at zero. A comparison of PID classifier score for data and simulation is shown on Fig. 2, which also indicates the cut on PID score $>$ 0.7 described in Sec. III.3.
Figure 1: Muon Classifier score comparison for $5\%$ of data and for predicted sample composition at the reconstruction level. The Monte Carlo normalization is scaled to match total event counts to data. The gray dashed line indicates the position of the cut on Muon Classifier score, with values to the left of the line removed from the final sample. Tinted bands around histogram lines represent $1\sigma$ statistical uncertainty.
Figure 2: PID score comparison for $5\%$ of data and for predicted sample composition at the reconstruction level. The Monte Carlo normalization is scaled to match total event counts to data. The gray dashed line indicates the position of the cut on PID score, with values to the left of the line removed from the final sample. Tinted bands around histogram lines represent $1\sigma$ statistical uncertainty.
Figure 3: Level 4 Noise Classifier score comparison for $5\%$ of data and for predicted sample composition at the reconstruction level. The Monte Carlo normalization is scaled to match total event counts to data. The gray dashed line indicates the position of the additional cut on the classifier score, with values to the left of the line removed from the final sample. Tinted bands around histogram lines represent $1\sigma$ statistical uncertainty.
@@
To further reduce non-neutrino background, the Muon Classifier score is required to be $>0.4$ for events to pass to the next step of selection. An additional cut of $>0.95$ is applied on the Level 4 Noise Classifier [11] to reduce remaining background. The performance of the Level 4 Noise Classifier score for our sample at the reconstruction level is shown in Fig. 3.
@@ III.2 Rejection of non-neutrino backgrounds
Another class of variables used to remove noise and low-quality events are calculated based on estimated photon speeds, under the assumption that each detected photon has started at the event interaction vertex and then traveled without scattering to the corresponding DOM. Photon speeds estimated in this way will not necessarily accurately represent the real event; however, the variables calculated using these are still useful, as they provide information on the event quality. In our event selection, we require the median estimated speed of all detected photons in an event to be $>0.4c$, where $c$ is the speed of light in vacuum. This helps us reject noise-induced events which may to have lower median speeds after being reconstructed with physics event hypothesis. We also require that in each event the fraction of detected photons with estimated speeds below zero or above $c$ is $<0.4$ because such pulses typically result from the PMT noise. Both of the thresholds were chosen based on Monte Carlo modeling. A photon speed below zero indicates that the pulse happened before the reconstructed interaction time, while photon speed above $c$ means that the pulse happened either too far from the reconstructed vertex or too early to be explained by a photon traveling from the vertex at the speed $c$. While a photon traveling in ice will have a speed of $c/n$, where $n$ is the refractive index of ice, the choice was made nevertheless to make a cut at the speed of $c$, to avoid penalizing events where a muon travels at speed higher than $c/n$ and then produces a Cherenkov photon near the DOM.
Restrictions are also applied to reject neutrino events with likely coincident atmospheric muons entering the detector. This is achieved by introducing cuts on events with pulses in the veto region. This region includes the top and the outermost optical modules of IceCube array, as these are more likely to detect light from atmospheric muons entering the detector. We limit the allowed number of pulses in the outermost layer of IceCube to be below 8. At the same time, we also require that there is no downward traveling trend in the pulse distribution in the top 15 layers of DOMs in IceCube, which corresponds to a vertical range of roughly 255 m.
We introduce vertex containment cuts on reconstructed vertex position to remove events with vertex located outside the DeepCore. We require that the vertical position of the vertex be in the range -500 $<z_{\mathrm{reco}}<$200 m from the vertical center of IceCube and that horizontal radial distance from the central string be $\rho_{\mathrm{reco}}<$ 300 m.
@@ III.3 Rejection of non-$\nu_{\mu}$ CC backgrounds
The remaining steps focus on selecting $\nu_{\mu}$ CC events with high-quality reconstruction. The primary event types targeted for the removal are very dim atmospheric muons, which leave few or no pulses on the veto region, and were therefore missed by previous selection steps, and non-$\nu_{\mu}$ CC neutrino interactions.
@@ III.4 Final level sample
A Monte Carlo estimate of the composition of the final sample is given in Table 1. Our selection results in a signal purity of 99%. The average ratio of neutrinos to antineutrinos is 2.4, which reflects the properties of the atmospheric neutrino flux, the neutrino cross section, and event selection efficiency dependence on true inelasticity. The data do not provide sufficient information to distinguish between neutrinos and antineutrinos, so the analysis is done for the sum of both.
The reconstruction resolutions for neutrino energy and inelasticity in the final simulated sample are presented in Fig. 4. The energy estimator shows a clear correlation with neutrino energy up to 500 GeV. Above that, the resolution degrades, because the typical size of the hadronic cascade approaches the size of DeepCore and a larger portion of the muon track could be located outside DeepCore. The inelasticity estimator also displays a correlation across all values, with deviations observed toward the edges, as is expected from one-sided errors in those regions. These two reconstructed quantities are used to extract the inelasticity distribution, as described in the next section.
@@ IV.1 General strategy
Figure 7: Fit of the parametrization (3) to inclusive inelasticity distributions for generator level genie Monte Carlo. Fits were done separately for neutrino and antineutrino inclusive inelasticity distributions in three energy bins, which were weighted to reflect energy distributions in energy bins at analysis level. Fit results for parameters $\langle y\rangle$ and $\log_{10}\lambda$ are shown in the bottom left of each plot.
Figure 8: Data/Monte Carlo agreement post-fit for the reconstructed inelasticity distribution in each of three $E_{\mathrm{reco}}$ bins.
This analysis is performed by passing data and simulation through the same selection and reconstruction steps, and then comparing their distributions in the space of observables chosen. The shape of the true inelasticity distribution of the simulation is parametrized, so it can be modified, along with nuisance parameters that describe the impact of different sources of uncertainties, in order to find the values that result in the best fit to the data. In this study, we compare a two-dimensional histogram in $E_{\mathrm{reco}}$ and $y_{\mathrm{reco}}$ with three bins in reconstructed energy between 100 GeV and 1 TeV and 11 equally sized bins in reconstructed inelasticity, between 0 and 1.
@@ IV.2 Parametrization of inelasticity
To parametrize the shape of the inelasticity distribution, our study takes advantage of the expected dependence of the neutrino-nucleon interaction cross section on inelasticity, as described in Sec. I. The functional form, used in the previous IceCube inelasticity study [7] and also described in Refs. [19, 20], is the probability distribution of inelasticity and depends on two parameters $\epsilon$ and $\lambda$,
@@
where $\sigma$ is total cross section for a given energy, $\frac{d\sigma}{dy}$ is the differential cross section as function of inelasticity for a given energy, and $N$ is a normalization factor. To treat the distribution as a probability density function it is normalized to 1, so
@@
It is important to emphasize that this parametrization is an effective description of the range of “reasonable” inelasticity distribution shapes; however, it has a number of limitations. To overcome these limitations, we use the $\log_{10}\lambda$ in our analysis and in the following sections. More details on this are provided in Appendix A.
The parametrization can be applied to both neutrino and antineutrino inelasticity distributions separately. Figure 7 shows Eq. (3) fitted to muon neutrino and antineutrino expectation from genie for three energy bins. However, due to our inability to differentiate between neutrino and antineutrino events, we use the following parametrization to describe the flux-averaged inelasticity distribution at energy $E_{\nu}$,
@@
where $\widetilde{\Phi}_{\nu(\bar{\nu})}$ is an energy-dependent fraction of neutrinos (antineutrinos) in our sample, which depends on atmospheric neutrino (antineutrino) flux $\Phi_{\nu(\bar{\nu})}$, the total neutrino-(antineutrino-)nucleon cross section $\sigma_{\nu(\bar{\nu})}$, and event selection, which is indicated as $\mathrm{sel.}$ in the equation. $\frac{dp}{dy}_{\nu(\bar{\nu})}$ denotes individual neutrino (antineutrino) inelasticity distributions. It is important to highlight that in Eq. (6) $E_{\nu}$ refers to both neutrino and antineutrino energy.